OpenStudy (anonymous):

f=1.5 +1.4(I-u) where I =0.025 and u=.044 please explain answer and process

OpenStudy (whpalmer4):

Just fill in the blanks: \[f = 1.5 + 1.4(0.025-0.044) =1.5+1.4(-0.019) = 1.5-0.0266 = 1.4734\]

OpenStudy (anonymous):

My answer key is showing the answer to be 8.4734 so I am confused

OpenStudy (anonymous):

I came up with same answer and solution as you, but key says is not correct

OpenStudy (whpalmer4):

There's a mistake somewhere, but it isn't in my arithmetic. Maybe the first number is supposed to be 8.5?

OpenStudy (anonymous):

what I came up with to match the answer key: .025 - .044 =.19 .019 x 1.4 = .0266 .0266 + 8.5 = 8.4734

OpenStudy (anonymous):

sorry - first line .019

OpenStudy (anonymous):

whpalmer4 - I am so sorry! It is 8.5 in the formula. I've been at these equations all week and ill to boot..... too many errors.... and first year algebra after no math for way too many years........... thank you for your help and attention to detail.

OpenStudy (whpalmer4):

"what I came up with to match the answer key: .025 - .044 =.19 .019 x 1.4 = .0266 .0266 + 8.5 = 8.4734" That should be: \[0.025-0.044 = -0.19\]\[-0.019*1.4=-0.0266\]\[-0.0266+8.5=8.4734\]The negative signs are important! Playing fast and loose with them will lead to mistakes.